OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

jimthompson5910 (jim_thompson5910):

if 1 + 3i is one root, then 1 - 3i is automatically one as well

jimthompson5910 (jim_thompson5910):

since all complex roots come in conjugate pairs

jimthompson5910 (jim_thompson5910):

so you really have 4 roots: 2, -4, 1 + 3i, 1-3i

jimthompson5910 (jim_thompson5910):

which means x = 2, x = -4, x = 1 + 3i, or x = 1 - 3i x - 2 = 0, x+4 = 0, x-1 = 3i or x - 1 = -3i x - 2 = 0, x+4 = 0, (x-1)^2 = (3i)^2 or (x - 1)^2 = (-3i) x - 2 = 0, x+4 = 0, (x-1)^2 = 9i^2 or (x - 1)^2 = 9i^2 x - 2 = 0, x+4 = 0, (x-1)^2 = 9i^2 x - 2 = 0, x+4 = 0, (x-1)^2 = 9(-1) x - 2 = 0, x+4 = 0, (x-1)^2 = -9 x - 2 = 0, x+4 = 0, (x-1)^2 + 9 = 0 I'll let you finish

OpenStudy (anonymous):

Thank You :)

jimthompson5910 (jim_thompson5910):

you're welcome